OpenStudy (anonymous):

how do u solve 3/2= 5/6+(-1/3) ??

OpenStudy (anonymous):

find a common denominator and multiply them each by that number to only have the numerators left. A common denominator among them would be 6, because 2 and 3 can both go into it. 6/2 = 3, so you multiply 3*3 6/6 = 1, so you multiply 5*1 6/3 = 2, so you multiply -1*2 you're left with \[9=5-2\] or 9=3, which doesn't really make sense at all. Is there more to the question?

OpenStudy (anonymous):

nope my books prob. wrong lol ill let my teacher know! thanks for trying to help!

OpenStudy (anonymous):

no problem ^^;